In numbers 1-3, find the length of A x B 1. A = 2i - 2j - k, B = i - k 2. A = 2i + 3j, B = -i + j 3. A = 2i - 2j + 4k, B = -i + j -2k In numbers 4-5, find a. A • B, |A|, |B| 4. A = 2i - 4j + √5k, B = -2i + 4j - √5k 5. A = (3/5)i + (4/5)k, B = 5i + 12j
1
Expert's answer
2021-10-28T08:52:55-0400

Cross product:

"1.\space\space A\times B=4i+3j+2k.\\
2.\space\space A\times B=5k.\\
3.\space\space A\times B=0."


Dot product:

"1.\space\space A·B=-25,\\
2.\space\space A·B=3."